The switch S in the circuit shown below has been o
Answer #1


after long time, the current through the C1 & C2 is zero

current in R1 , R2 = V/(R1+R2)


I = 12/(110+220)


I = 0.036 A

potential across R2 = V2 = I*R2 = 0.036*220 = 7.92 V

Q = Cs*V = (C1*c2)*V/(C1+C2)

Q = (40*10^-6*80*10^-6)*7.92/(40*10^-6+80*10^-6)

Q = 211.2*10^-6 C